### v1

# ForLion: An R Package for Finding Optimal Experimental Designs with Mixed Factors

Abstract
Optimal design is crucial for experimenters to maximize the information collected from experiments and estimate the model parameters most accurately. ForLion algorithms have been proposed to find D-optimal designs for experiments with mixed types of factors. In this paper, we introduce the ForLion package which implements the ForLion algorithm to construct locally D-optimal designs and the EW ForLion algorithm to generate robust EW D-optimal designs. The package supports experiments under linear models (LM), generalized linear models (GLM), and multinomial logistic models (MLM) with continuous, discrete, or mixed-type factors. It provides both optimal approximate designs and an efficient function converting approximate designs into exact designs with integer-valued allocations of experimental units. Tutorials are included to show the package’s usage across different scenarios.
1 Introduction
The study of optimal designs can be traced back to Smith (1918) on regression problems for univariate polynomials of order up to six (Fedorov and Leonov, 2014). Later in the 20th century, the optimal design theory has been expanded to encompass various statistical models and optimality criteria (Fedorov, 1972; Silvey, 1980; Pukelsheim, 1993; Atkinson et al., 2007; Fedorov and Leonov, 2014). Despite significant advancements on linear regression models, generalized linear models (GLM) for more general univariate responses (McCullagh and Nelder, 1989; Dobson and Barnett, 2018; Khuri et al., 2006; Stufken and Yang, 2012) and multinomial logistic models (MLM) for categorical responses (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020) have been widely used in practice, but are much more difficult in the optimal design theory, because their Fisher information matrices depend on model parameters.
Package AlgDesign (version 1.2.1.2, see Wheeler and Braun (2025)) offers tools to construct optimal designs with exact and approximate allocations, focusing on linear models including polynomial forms. It allows mixed factors by generating a finite candidate list of experimental settings and employs the Fedorov’s exchange algorithm (Fedorov, 1972) for optimization under D-, A-, and I-criteria. Package OptimalDesign (version 1.0.2.1, see Harman and Filová (2025); Harman et al. (2020)) finds D-, A-, I-, and c-efficient designs for linear models (LM), generalized linear models (GLM), some dose-response, and certain survival models, with mixed factors by discretizing the continuous factors first. Package ICAOD (version 1.0.1, see Masoudi et al. (2020)) provides tools for finding locally, minimax, and Bayesian D-optimal designs, as well as user-specified optimality criteria, for nonlinear statistical models. It implements the imperialist competitive algorithm (Masoudi et al., 2022) for designs involving continuous factors only. Package PFIM (version 6.1, see Mentré et al. (2024); Dumont et al. (2018)) computes D-optimal designs for nonlinear mixed-effects models (NLMEM), which is particularly useful for pharmacokinetic/pharmacodynamic models (PKPD). It handles designs either with discrete factors only or with continuous factors only. Package idefix (version 1.1.0, see Traets et al. (2025, 2020)) is designed to generate D-efficient and Bayesian D-efficient optimal designs for multinomial logit (MNL) and mixed logit (MIXL) models for discrete choice experiments (DCE). However, these existing packages are designed for experiments with discrete factors only, continuous factors only, or mixed factors by discretizing the continuous factors first. There has been limited progress in constructing efficient designs that incorporate both discrete and continuous factors (Huang et al., 2024).
The ForLion package, available at the Comprehensive R Archive Network (CRAN, https://CRAN.R-project.org/package=ForLion), addresses this gap by providing a solution for experimental designs involving both types of factors under fairly general parametric statistical models using the ForLion algorithm proposed by Huang et al. (2024). Different from discretizing the continuous factors first, it conducts an exhaustive search within a mixed-factor design space, and minimizes the number of distinct experimental settings while maintaining high efficiency of designs. It covers linear models (LM), generalized linear models (GLM), and general multinomial logistic models (MLM). Note that the MLM here is much broader than the MNL for discrete choice experiments. It includes three more classes of models for ordinal responses, namely cumulative, adjacent-categories, and continuation-ratio logit models (Bu et al., 2020). Furthermore, to overcome the issue caused by the dependence of Fisher information matrices on model parameters, we also implement the EW ForLion algorithm proposed by Lin et al. (2025) for constructing robust optimal designs against unknown model parameters for LM, GLM, and MLM as well.
In Section 2, we outline the theoretical foundations of the ForLion package, including the ForLion algorithm, the Lift-one algorithm, the EW ForLion algorithm, and the rounding algorithm. In Section 3, we introduce the structure and function arguments of the ForLion package, followed by illustrative examples of the package applications, as well as the interpretations of the key results in Section 4. We summarize and conclude in Section 5.
2 Method
We consider a mixed-factors experiment under a general statistical model $M(\mathbf{x};\boldsymbol{\theta})$ with parameter(s) $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$ and experimental setting ${\mathbf{x}}\in{\cal X}\subset\mathbb{R}^{d}$, where $\boldsymbol{\Theta}$ is called the parameter space, and ${\cal X}$ is called the design region or design space. For typical applications, ${\cal X}$ is compact. Among the $d$ factors, without loss of generality, we assume that the first $k$ factors are continuous and the last $d-k$ are discrete, for $0\leq k\leq d$. Following Lin et al. (2025), in package ForLion, we cover three scenarios: (i) if $k=0$, ${\cal X}={\cal D}$ consists of a predetermined finite list of level combinations of discrete factors; (ii) if $11\leq k\leq d-$ 1, ${\cal X}=\prod_{j=1}^{k}I_{j}\times{\cal D}$, with $I_{j}=[a_{j},b_{j}]$ being a finite closed interval; and (iii) if $k=d$, ${\cal X}=\prod_{j=1}^{d}I_{j}$
An experimental design $\boldsymbol{\xi}$ considered in this paper consists of $m$ distinct design points, $\mathbf{x}_{1},\dots,\mathbf{x}_{m}$ $\in{\cal X}$, and real-valued proportions ${\mathbf{w}}=(w_{1},\ldots,w_{m})^{\top}$ satisfying $w_{i}\geq 0$ for each $i$ and $\sum_{i=1}^{m}w_{i}=1$, known as an approximate allocation of the experimental units. In practice, we also look for integer-valued assignments ${\mathbf{n}}=(n_{1},\ldots,n_{m})^{\top}$ given $N=\sum_{i=1}^{m}n_{i}$ , known as an exact allocation. The collection of approximate designs is denoted by $\boldsymbol{\Xi}=\{\boldsymbol{\xi}=\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m\}\mid m\geq 1,{\mathbf{x}}_{i}\in{\cal X},w_{i}\geq 0,\sum_{i=1}^{m}w_{i}=1\}$. Under regularity conditions, the Fisher information matrix associated with the design $\boldsymbol{\xi}$ can be denoted as ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})=\sum_{i=1}^{m}w_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})\in\mathbb{R}^{p\times p}$, where ${\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$ is the Fisher information associated with ${\mathbf{x}}_{i}$
When the experimenter has a good idea about the values of $\boldsymbol{\theta}$, following Huang et al. (2024), we look for a locally D-optimal design that maximizes $f_{\boldsymbol{\theta}}(\boldsymbol{\xi})=|\mathbf{F}(\boldsymbol{\xi},\boldsymbol{\theta})|$ by implementing the ForLion algorithm (see Section 2.1). In practice, however, the experimenter may not be certain about the true parameter values. In package ForLion, we offer two options for the users. With a prespecified prior distribution or probability measure $Q(\cdot)$ on $\boldsymbol{\Theta}$, we look for an EW D-optimal design (Atkinson et al., 2007; Yang et al., 2016, 2017; Bu et al., 2020; Huang et al., 2025b; Lin et al., 2025) that maximizes
$f_{\rm EW}(\boldsymbol{\xi})=|E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|=\left|\int_{\boldsymbol{\Theta}}{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})Q(d\boldsymbol{\theta})\right|$ (1)
by implementing the EW ForLion algorithm proposed by Lin et al. (2025), also called an integral-based EW D-optimal design. When a dataset from a pilot study is available, or if the integrals in (1) are difficult to calculate, we look for a sample-based EW D-optimal design (Lin et al., 2025) that maximizes
$f_{\rm SEW}(\boldsymbol{\xi})=|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|=\left|\frac{1}{B}\sum_{j=1}^{B}{\mathbf{F}}(\boldsymbol{\xi},\hat{\boldsymbol{\theta}}_{j})\right|\ ,$ (2)
where $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ are either estimated from bootstrapped samples from the pilot dataset, or sampled from the prior distribution $Q(\cdot)$ on $\boldsymbol{\Theta}$ (see Section 2.3).
According to Huang et al. (2024), unlike stochastic optimization algorithms such as particle swarm optimization (PSO), the designs found by the ForLion algorithm are guaranteed to be locally D-optimal. Based on Lin et al. (2025), the designs constructed by the EW ForLion algorithm are guaranteed to be EW D-optimal designs, which are not only robust against unknown parameters, but also computationally much easier to find than the traditional Bayesian D-optimal designs, while keeping high-efficiency (Yang et al., 2016, 2017; Bu et al., 2020; Lin et al., 2025).
2.1 ForLion algorithm
In this section, we assume that $\boldsymbol{\theta}\in\boldsymbol{\Theta}$ is known. To simplify notations, we let ${\mathbf{F}}_{\mathbf{x}}$ and ${\mathbf{F}}(\boldsymbol{\xi})$ stand for ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})$ and ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})$, respectively. The ForLion algorithm looks for a locally D-optimal design $\boldsymbol{\xi}_{*}\in\boldsymbol{\Xi}$ that maximizes $|{\mathbf{F}}(\boldsymbol{\xi})|$. It begins with an initial design $\boldsymbol{\xi}_{0}\in\boldsymbol{\Xi}$ in Step $1^{\circ}$ (see Figure 1) satisfying $|{\mathbf{F}}({\boldsymbol{\xi}}_{0})|>0$, with a prespecified distance threshold $\delta_{0}>0$ between any two design points, such that, $\|{\mathbf{x}}_{i}^{(0)}-{\mathbf{x}}_{j}^{(0)}\|\geq\delta_{0}$ for any $i\neq j$. To minimize the number of distinct experimental settings, which in practice often indicates reduced experimental time and cost, at the beginning of each round of iterations, the ForLion algorithm reduces the number of design points by merging those points in close distance (i.e., less than $\delta$), which is performed only if the resulting design still satisfies $|{\mathbf{F}}({\boldsymbol{\xi}}_{\rm merge})|>0$ (see Step $2^{\circ}$ in Figure 1). In Step $3^{\circ}$, the lift-one algorithm (Yang and Mandal, 2015; Yang et al., 2016; Huang et al., 2025b) is employed to update the approximate allocation for the current design $\boldsymbol{\xi}_{t}$ In Step $4^{\circ}$, the design points with zero allocation are removed from the design.
In Step $5^{\circ}$, the ForLion algorithm identifies a new design point $\mathbf{x}^{*}$ that maximizes the sensitivity function $d(\mathbf{x},\boldsymbol{\xi}_{t})=\operatorname{tr}(\mathbf{F}(\boldsymbol{\xi}_{t})^{-1}\mathbf{F}_{\mathbf{x}})$. According to Theorem 2.2 of Fedorov and Leonov (2014) and Theorem 1 in Huang et al. (2024), if $d(\mathbf{x}^{*},\boldsymbol{\xi}_{t})\leq p$, then $\boldsymbol{\xi}_{t}$ is D-optimal; otherwise, $\mathbf{x}^{*}$ is added to $\boldsymbol{\xi}_{t}$ with an initial zero allocation in Step $6^{\circ}$, and the process returns to Step $2^{\circ}$. This iterative process continues until the condition $d(\mathbf{x},\boldsymbol{\xi}_{t})\leq p$ is met, at which the design is reported as the D-optimal design. The outline of the ForLion algorithm is provided in Figure 1, with detailed descriptions available in Algorithm 1 of Huang et al. (2024).
Figure 1: An outline of the ForLion algorithm under a general statistical model.
To speed up the ForLion algorithm for LM and GLM, we replace Steps $1^{\circ}$, $3^{\circ}$, $5^{\circ}$, and $6^{\circ}$ in Figure 1 with Steps $1^{\prime}$, $3^{\prime}$, $5^{\prime}$, and $6^{\prime}$ in Figure 2, respectively, and implement the analytical solutions for GLM derived by Yang and Mandal (2015) and Huang et al. (2024). More specifically, in Step $1^{\prime}$, we replace the random initial design with a minimally supported uniform design. In Step $3^{\prime}$, we adopt the analytic solutions provided by Yang and Mandal (2015) for the lift-one algorithm (see Section 2.2) under a GLM. In Step $5^{\prime}$, we adopt the simplified form of the sensitivity function as described in Theorem 4 of Huang et al. (2024). In Step $6^{\prime}$, we assign an initial weight $\alpha_{t}$ instead of zero to the new design point $\mathbf{x}^{*}$, based on Theorem 5 of Huang et al. (2024). Those modifications accelerate the ForLion algorithm for LM and GLM. The specialized ForLion algorithm is illustrated by Figure 2, with further details provided in Section 4 of Huang et al. (2024).
Figure 2: An outline of the GLM-adapted ForLion algorithm.
2.2 Lift-one algorithm
The lift-one algorithm implemented in Step $3^{\circ}$ of the ForLion algorithm was originally proposed by Yang et al. (2016) and Yang and Mandal (2015) for GLMs, and then extended for cumulative link models by Yang et al. (2017), and MLM by Bu et al. (2020). A general version and a constrained version of it can be found in Huang et al. (2025b) and their Supplementary Material. It is highly efficient for finding the D-optimal approximate allocation ${\mathbf{w}}=(w_{1},\ldots,w_{m})^{\top}$ for a given set of design points $\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{m}\}$. By adjusting the $i$th weight $w_{i}<1$ while rescaling the remaining weights proportionally, the adjusted allocation is given by
$\mathbf{w}_{i}(z)=\left(\frac{1-z}{1-w_{i}}w_{1},\dots,\frac{1-z}{1-w_{i}}w_{i-1},z,\frac{1-z}{1-w_{i}}w_{i+1},\dots,\frac{1-z}{1-w_{i}}w_{m}\right)^{\top}$
with $z\in[0,1]$, which converts a multi-dimensional optimization problem to a one-dimensional optimization problem. By adjusting ${\mathbf{w}}_{i}(z)$ for each $i=1,\dots,m$ in a random order, the converged allocation is guaranteed to be D-optimal.
According to a comprehensive comparison study by Yang et al. (2016), the lift-one algorithm is much faster than conventional optimization techniques, including Nelder-Mead, quasi-Newton, and simulated annealing, as well as popular design algorithms such as Fedorov-Wynn, multiplicative, and cocktail algorithms (see Huang et al. (2024) for a good review). It often achieves a design with a reduced number of support points, that is, the design points with positive weights.
2.3 EW ForLion algorithm
In this section, the model parameter vector $\boldsymbol{\theta}\in\boldsymbol{\Theta}$ is assumed to be unknown. Instead, either a prior distribution $Q(\cdot)$ on $\boldsymbol{\Theta}$ or a dataset obtained from a previous study is available.
Given a prior distribution or probability measure $Q(\cdot)$ on $\boldsymbol{\Theta}$, we adopt the EW ForLion algorithm proposed by Lin et al. (2025) to find an intergral-based EW D-optimal design $\boldsymbol{\xi}_{*}\in\boldsymbol{\Xi}$ that maximizes $f_{\rm EW}(\boldsymbol{\xi})=|\int_{\boldsymbol{\Theta}}{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})Q(d\boldsymbol{\theta})|$ as defined in (1). Different from the ForLion algorithm maximizing $f_{\boldsymbol{\theta}}(\boldsymbol{\xi})=|\mathbf{F}(\boldsymbol{\xi},\boldsymbol{\theta})|$ with a prespecified $\boldsymbol{\theta}$, the EW ForLion algorithm targets the expectation of the Fisher information matrix, $E\left\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\right\}$. By computing the entry-wise expectation with respect to the prior distribution $Q(\cdot)$ on $\boldsymbol{\Theta}$, we obtain a $p\times p$ matrix as well. Commonly used prior distributions include uniform priors on bounded intervals, normal priors on $\mathbb{R}$, and Gamma priors on $(0,\infty)$ (see, e.g., Huang et al. (2025b)). In this package, we calculate the corresponding integrals by using the hcubature() function in package cubature (Narasimhan et al., 2025), which applies an adaptive multidimensional integration method by subdividing hyper-rectangular domains.
Alternatively, if a dataset from a previous or pilot study is available, we may bootstrap it for $B$ times (e.g., $B=1000$). For the $j$th bootstrapped dataset, we fit the statistical model and obtain the parameter estimates $\hat{\boldsymbol{\theta}}_{j}$ , for $j=1,\ldots,B$. In this package, we implement the EW ForLion algorithm to find a sample-based EW D-optimal design $\boldsymbol{\xi}_{*}$ , which maximizes $f_{\rm SEW}(\boldsymbol{\xi})=|\sum_{j=1}^{B}{\mathbf{F}}(\boldsymbol{\xi},\hat{\boldsymbol{\theta}}_{j})|$ as defined in (2). In practice, if the expected Fisher information matrix is difficult to calculate with respect to $Q(\cdot)$, e.g., for some MLM (Lin et al., 2025), we may also simulate $\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}$ from $Q(\cdot)$, and look for a sample-based EW D-optimal design. According to Lin et al. (2025), the resulting designs are fairly robust against different sets of parameter vectors in terms of relative efficiency.
When there is no confusion, we let ${\mathbf{F}}_{\mathbf{x}}$ represent $\int_{\boldsymbol{\Theta}}{\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ for integral-based EW optimality, or $B^{-1}\sum_{j=1}^{B}{\mathbf{F}}({\mathbf{x}},\hat{\boldsymbol{\theta}}_{j})$ for sample-based EW optimality. Similarly, we let ${\mathbf{F}}(\boldsymbol{\xi})$ represent $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for integral-based EW optimality, or $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ for sample-based EW optimality. Then Figure 1 may also be used for illustrating the EW ForLion algorithm, with more detailed descriptions in Algorithm 1 of Lin et al. (2025).
2.4 Rounding algorithm
Both the ForLion and EW ForLion algorithms are able to find optimal experimental settings in a continuous or mixed design region. In practice, the suggested experimental settings may need to be rounded up due to the sensitivity level of the experimental device or environmental control. For example, 149.2116 Gy as the gamma radiation level may need to be rounded up to 149.2 Gy due to the sensitivity level of the radiation device (see the emergence of house flies example in Section 4.1). On the other hand, the obtained optimal approximate design may also need to be converted to an exact design given a total number $N$ of experimental units.
In this package, we adopt the rounding algorithm proposed by Lin et al. (2025) to convert an approximate design obtained by the ForLion or EW ForLion algorithms on a continuous or mixed region to an exact design with user-specified levels of grid points and $N$. It starts by merging design points based on a specified distance measure and a merging threshold $\delta_{2}$ (see Step $1^{\circ}$ in Figure 3). Then the levels of the continuous factors are rounded to the nearest multiples of user-defined grid levels (see Step $2^{\circ}$). As for the approximate allocation $w_{i}$ , the rounding algorithm initializes the corresponding integer allocation $n_{i}=\lfloor{N\times w_{i}}\rfloor$, the largest integer no more than $Nw_{i}$ , and then allocates the remaining experimental units one by one to maximize the objective function (see Step $3^{\circ}$). The resulting exact design is reported in Step $4^{\circ}$. The outline of the rounding algorithm is displayed in Figure 3, with more details provided in Algorithm 2 of Lin et al. (2025) .
Compared with optimal designs constructed directly on the same set of grid points, the design rounded from a ForLion design costs much less time, contains less distinct experimental settings, and maintains a high relative efficiency with respect to the ForLion design (Lin et al., 2025).
Figure 3: An outline of the rounding algorithm.
3 ForLion Package Structure
The current version (0.3.0) of the ForLion package (Huang et al., 2025a) supports finding D-optimal designs for experiments under parametric models with discrete factors only, continuous factors only, or mixed factors, including linear models (LM or GLM with “identity” link), logistic models (GLM with “logit” link) and other GLMs for binary responses (GLM with “probit”, “cloglog”, “loglog”, and “cauchit” links), Poisson models (GLM with “log” link), baseline-category logit models or multiclass logistic models (MLM with “baseline” link), cumulative logit models (MLM with “cumulative” link), adjacent-categories logit models (MLM with “adjacent” link), and continuation-ratio logit models (MLM with “continuation” link). Its key functions and structure are displayed in Figure 4.
Figure 4: The key functions in the ForLion package and the corresponding structure.
More specifically, to construct locally D-optimal designs, ForLion provides two key functions by implementing the ForLion algorithm (Huang et al., 2024):
  • ForLion_MLM_Optimal for finding locally D-optimal approximate designs under an MLM.
  • ForLion_GLM_Optimal for finding locally D-optimal approximate designs under a GLM, which covers LM with “identity” link as a special case.
The specialized arguments and their descriptions for these functions are summarized in Table 1.
To construct robust D-optimal designs against unknown parameter values, ForLion package offers two key functions as well by implementing the EW ForLion algorithm (Lin et al., 2025):
  • EW_ForLion_MLM_Optimal for finding EW D-optimal approximate designs under an MLM.
  • EW_ForLion_GLM_Optimal for finding EW D-optimal approximate designs under a GLM.
Details about additional arguments for implementing these functions are listed in Table 2.
To obtain exact designs with user-defined grid points and number of experimental units from the D-optimal approximate designs, ForLion package provides the following two functions by implementing the rounding algorithm proposed by Lin et al. (2025):
  • MLM_Exact_Design for obtaining exact designs under an MLM.
  • GLM_Exact_Design for obtaining exact designs under a GLM.
Finally, a list of arguments commonly used in all the key functions is presented in Table 3.
When finding robust designs under an MLM, such as a cumulative logit model, the feasible parameter space may not be rectangular (Bu et al., 2020; Lin et al., 2025), and the computation for integral-based EW D-optimality (see (1)) is much more difficult, especially with a moderate or large number $J$ of response categories. In ForLion package, we adopt the sample-based EW D-optimality (see (2)) for MLMs. As for GLMs, we allow users to choose integral-based or sample-based D-optimality for deriving robust approximate and exact designs (see also Section 2.3).
Table 1: Special arguments for two ForLion functions
Function Argument Description
ForLion_MLM_Optimal J Number of response categories in an MLM.
h.prime Function to obtain $\partial{\mathbf{X}}_{\mathbf{x}}/\partial{\mathbf{x}}$.
bvec Vector of assumed parameter values.
Fi.func Function to calculate row-wise Fisher information Fi, default Fi_MLM_func.
delta0 Merging threshold for initial design, such that, $\|{\mathbf{x}}_{i}^{(0)}-{\mathbf{x}}_{j}^{(0)}\|\geq\delta_{0}$ for $i\neq j$, default $10^{-5}$.
epsilon Tuning parameter as converging threshold, such that, a nonnegative number is regarded as numerical zero if less than epsilon, default $10^{-12}$.
random TRUE or FALSE, whether or not to repeat the lift-one step multiple times with random initial allocations, default FALSE.
nram Number of times repeating the lift-one step with random initial allocations, valid only if random is TRUE, default 3.
random.initial TRUE or FALSE, whether or not to repeat the whole procedure multiple times with random initial designs, default FALSE.
nram.initial Number of times repeating the whole procedure with random initial designs, valid only if random.initial is TRUE, default 3.
optim_grad TRUE or FALSE, default FALSE, whether to use analytical or numerical gradient function when searching for a new design point.
ForLion_GLM_Optimal logscale TRUE or FALSE, whether or not to run the lift-one step in log-scale, i.e., using liftoneDoptimal_log_GLM_func() or liftoneDoptimal_GLM_func().
Table 2: Special arguments for two EW_ForLion functions
Function Argument Description
EW_ForLion_MLM_Optimal bvec_matrix Matrix of bootstrapped or simulated parameter values.
EW_Fi.func Function to calculate entry-wise expectation of Fisher information Fi, default EW_Fi_MLM_func.
optim_grad TRUE or FALSE, default FALSE, whether to use analytical or numerical gradient function when searching for a new design point.
EW_ForLion_GLM_Optimal joint_Func_b Prior distribution function of model parameters
Integral_based TRUE or FALSE, whether or not integral-based EW D-optimality is used, FALSE indicates sample-based EW D-optimality is used.
b_matrix Matrix of bootstrapped or simulated parameter values.
Lowerbounds Vector of lower ends of ranges of prior distribution for model parameters.
Upperbounds Vector of upper ends of ranges of prior distribution for model parameters.
logscale TRUE or FALSE, whether or not to run the lift-one step in log-scale, i.e., using EW_liftoneDoptimal_log_GLM_func() or EW_liftoneDoptimal_GLM_func().
Table 3: Common arguments in four key functions
Argument Description
xlist_fix List of discrete-factor experimental settings under consideration, default NULL indicating a list of all possible discrete-factor experimental settings will be used.
n.factor Vector of numbers of distinct levels, “0” indicating continuous factors that always come first, “2” or more for discrete factors, and “1” not allowed.
factor.level List of distinct factor levels, “(min, max)” for continuous factors that always come first, finite sets for discrete factors.
var_names Names for the design factors, Must have the same length as factor.level; default “X1”, “X2”, “X3”, etc.
hfunc Function for generating the corresponding model matrix or predictor vector, given an experimental setting or design point.
link Link function, for GLMs including “logit” (default), “probit”, “cloglog”, “loglog”, “cauchit”, “log”, and “identity” (LM); for MLMs including “continuation” (default), “baseline”, “cumulative”, and “adjacent”.
reltol Relative tolerance as converging threshold, default $10^{-5}$.
delta Relative difference as merging threshold for the merging step, two points satisfying $\|{\mathbf{x}}_{i}^{(t)}-{\mathbf{x}}_{j}^{(t)}\|<\delta$ may be merged, default 00, can be different from delta0 for the initial design.
maxit Maximum number of iterations allowed, default $100$.
rowmax Maximum number of points in initial designs, default NULL indicating no restriction.
Xini Initial list of design points, default NULL indicating automatically generating an initial list of design points.
4 Examples
In this section, we demonstrate by examples how to use the ForLion package to obtain locally D-optimal approximate designs, EW D-optimal approximate designs according to integral-based or sample-based EW D-optimality, and the corresponding exact designs, under different scenarios.
4.1 An MLM example: Emergence of house flies experiment
We demonstrate the implementation of the ForLion_MLM_Optimal() function by exploring an emergence of house flies experiment. The original experiment, described by Itepan (1995), involved $N=3500$ pupae uniformly assigned to $m=7$ different levels of a gamma radiation device (in Gy): $x_{i}=80,100,120,140,160,180,200$. The study recorded three categorical outcomes, namely unopened, opened but died, and opened and emerged, which clearly have an order or structure. A continuation-ratio non-proportional odds (npo) model with $J=3$, as a special case of MLM, was considered in the literature (Zocchi and Atkinson, 1999; Bu et al., 2020; Ai et al., 2023):
$\log\left(\frac{\pi_{i1}}{\pi_{i2}+\pi_{i3}}\right)$ $=$ $\beta_{11}+\beta_{12}x_{i}+\beta_{13}x_{i}^{2}\ ,$
$\log\left(\frac{\pi_{i2}}{\pi_{i3}}\right)$ $=$ $\beta_{21}+\beta_{22}x_{i}\ ,$
with parameters $\hat{\boldsymbol{\theta}}=(\beta_{11},\beta_{12},\beta_{13},\beta_{21},\beta_{22})^{\top}=(-1.935,-0.02642,0.0003174,-9.159,$ $0.06386)^{\top}$ fitted from the pilot study by Itepan (1995), where $i=1,\ldots,m$.
Following Ai et al. (2023) and Huang et al. (2024), we reconsider the experiment with a continuous range for the gamma radiation levels, namely $x_{i}\in\mathcal{X}=[0,200]$.
Finding a locally D-optimal approximate design
We start by defining the assumed model parameter values theta, the design matrix function hfunc.temp, along with its gradient function hprime.temp:
> theta = c(-1.935, -0.02642, 0.0003174, -9.159, 0.06386)
> hfunc.temp = function(y){
+     matrix(data=c(1,y,y*y,0,0,0,0,0,1,y,0,0,0,0,0), nrow=3,
      ncol=5, byrow=TRUE)} #forlion hfunc: design matrix
> hprime.temp = function(y){
+ list(matrix_1 =matrix(data=c(0, 1, 2*y, 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0),
+ nrow=3, ncol=5, byrow=TRUE))} #forlion h.prime: gradient of hfunc
Next, the ForLion_MLM_Optimal() function can be applied to find a locally D-optimal approximate design under the continuation-ratio npo model using the following R codes:
> set.seed(123)
> forlion.design = ForLion_MLM_Optimal(J=3, n.factor=c(0),
+ factor.level=list(c(0, 200)), hfunc=hfunc.temp, h.prime=hprime.temp,
+ bvec=theta, link="continuation", Fi.func=Fi_MLM_func, delta0=1e-6,
+ epsilon=1e-12, reltol=1e-8, delta=0.15, maxit=1000, random=TRUE,
+ nram=3, rowmax=NULL, Xini=NULL, random.initial=TRUE, nram.initial=3)
For the function arguments, we specify n.factor=c(0) to indicate that there is no discrete factor in this experiment. The range of the only continuous factor is defined through factor.level=list(c(0, 200)), which is a closed interval. The obtained approximate design by the ForLion algorithm can be summarized and output using print() function as below:
> print(forlion.design)
Design Output
Count  X1        Allocation
1      103.5300  0.3981
2        0.0000  0.2027
3      149.2116  0.3992
m:
[1] 3
det:
[1] 54016299
convergence:
[1] TRUE
min.diff:
[1] 45.6816
x.close:
[1] 103.5300 149.2116
itmax:
[1] 23
The above Design Output identifies three design points $x_{1}=0.0000$, $x_{2}=103.5300$, and $x_{3}=149.2116$, along with the approximate allocations $w_{1}=0.2027$, $w_{2}=0.3981$, and $w_{3}=0.3992$. Here m=3 matches the number of distinct design points. The determinant det of the Fisher information matrix associated with the obtained design $\boldsymbol{\xi}=\{(x_{i},w_{i})\mid i=1,2,3\}$ is $\left|\mathbf{F}(\boldsymbol{\xi},\hat{\boldsymbol{\theta}})\right|=54,016,299$. Furthermore, convergence: TRUE indicates that the algorithm converged successfully. The minimum Euclidean distance min.diff among the distinct design points is $45.6816$, and the closest pair x.close of the design points is $103.5300$ and $149.2116$. Finally, itmax=23 indicates that the number of iterations spent by the algorithm is $23$.
Obtaining exact designs from the locally D-optimal approximate design $\boldsymbol{\xi}$
Among the three reported design points in $\boldsymbol{\xi}$, two of them, namely $x_{2}=103.5300$ and $x_{3}=149.2116$ in Gy, may not be feasible in practice due to the sensitivity level of the gamma radiation device. For example, if the device can only allow the gamma ratiation level to be set as a multiple of 0.1 Gy, we may use the MLM_Exact_Design() function with grid level $L=0.1$ as follows:
ForLion_exact_L0.1<-MLM_Exact_Design(J=3, k.continuous=1,design_x=
+ forlion.design\$x.factor,design_p=forlion.design\$p,det.design=forlion.design\$det,
+ p=5,ForLion=TRUE,bvec=theta,delta2=1,L=0.1,N=3500,hfunc=hfunc.temp,
+ link="continuation")
where J=3 represents the number of response categories, k.continuous=1 indicates that there is only one continuous variable, namely the gamma radiation level, design_x is the set of design points in $\boldsymbol{\xi}$, design_p stores the corresponding approximate allocations, det.design is the maximized determinant of the Fisher information matrix, p=5 stands for the number of parameters, and N=3500 is the total number of experimental units. Furthermore, The argument ForLion=TRUE indicates that the approximate design is obtained by the ForLion algorithm, while ForLion=FALSE corresponds to the EW ForLion algorithm. The converted exact design, denoted by $\boldsymbol{\xi}_{\rm exact}$ is output as below:
> print(ForLion_exact_L0.1)
Design Output
Count  X1        Allocation
1      103.5000  0.3981
2        0.0000  0.2027
3      149.2000  0.3992
ni.design:
[1] 1393  710 1397
rel.efficiency:
[1] 0.9999989
The output of the exact design $\boldsymbol{\xi}_{\rm exact}$ shows that it contains three design points, namely $x_{1}=0$, $x_{2}=103.5$, $x_{3}=149.2$, along with the corresponding integer-valued allocations ni.design, namely $n_{1}=710$, $n_{2}=1393$, $n_{3}=1397$. The relative efficiency rel.efficiency of $\boldsymbol{\xi}_{\rm exact}$ with respect to the D-optimal approximate design $\boldsymbol{\xi}$, is $(|{\mathbf{F}}(\boldsymbol{\xi}_{\rm exact},\hat{\boldsymbol{\theta}})|/|{\mathbf{F}}(\boldsymbol{\xi},\hat{\boldsymbol{\theta}})|)^{1/p}=0.9999989$, or $99.99989\%$, which is highly efficient.
For illustration purposes, we repeat the procedure for some other grid levels, namely $L=1,5,10$, and $20$ as well. The corresponding exact designs, as well as $L=0.1$, are summarized in Table 4. According to Table 4, the relative efficiency decreases as the rounding level $L$ increases. In practice, the experimenters may choose an appropriate $L$ as a compromise between the experimental requirements and the relative efficiency.
Table 4: Exact designs for house files experiment with different rounding levels and $N=3500$
$i$ L=0.1 L=1 L=5 L=10 L=20
$x_{i}$ $n_{i}$ $x_{i}$ $n_{i}$ $x_{i}$ $n_{i}$ $x_{i}$ $n_{i}$ $x_{i}$ $n_{i}$
1 0 710 0 710 0 710 0 710 0 710
2 103.5 1393 104 1393 105 1393 100 1393 100 1393
3 149.2 1397 149 1397 150 1397 150 1397 140 1397
Rel. Efficiency 99.99989% 99.98448% 99.93424% 99.48902% 94.65724%
Finding a sample-based EW D-optimal approximate design
Next, we look for an EW D-optimal approximate design rather than a locally D-optimal one with the prespecified $\hat{\boldsymbol{\theta}}$. It is more robust against misspecified parameter values (Lin et al., 2025). More specifically, we first draw $B=1,000$ bootstrapped samples from the previous data (Table 1 in Zocchi and Atkinson (1999)), and then obtain the corresponding parameter estimates $\hat{\boldsymbol{\theta}}_{j}$ from the $j$th bootstrapped sample, for $j=1,\ldots,B$. The $1,000$ parameter vectors are stored as matrix theta_matrix as below:
> n=1000
> set.seed(2024)
> tt1=rmultinom(n, size = 500, prob = c(62/500, 5/500, 433/500))
> tt1t=t(tt1)
> tt1x=cbind(80,6400,tt1t)
> tt2=rmultinom(n, size = 500, prob = c(94/500, 24/500, 382/500))
> tt2t=t(tt2)
> tt2x=cbind(100,10000,tt2t)
> tt3=rmultinom(n, size = 500, prob = c(179/500, 60/500,  261/500))
> tt3t=t(tt3)
> tt3x=cbind(120,14400,tt3t)
> tt4=rmultinom(n, size = 500, prob = c(335/500, 80/500, 85/500))
> tt4t=t(tt4)
> tt4x=cbind(140,19600,tt4t)
> tt5=rmultinom(n, size = 500, prob = c(432/500, 46/500,  22/500))
> tt5t=t(tt5)
> tt5x=cbind(160,25600,tt5t)
> tt6=rmultinom(n, size = 500, prob = c( 487/500, 11/500,  2/500))
> tt6t=t(tt6)
> tt6x=cbind(180,32400,tt6t)
> tt7=rmultinom(n, size = 500, prob = c(498/500, 2/500, 0/500))
> tt7t=t(tt7)
> tt7x=cbind(200,40000,tt7t)
> theta_matrix=rep(0,n*5); dim(theta_matrix)=c(n,5)
> for (i in 1:n) {
+ data = rbind(tt1x[i,], tt2x[i,], tt3x[i,], tt4x[i,], tt5x[i,], tt6x[i,],
+ tt7x[i,])
+ colnames(data)=c(’x1’,’x2’,’y1’,’y2’,’y3’)
+  data1=data.frame(data)
+  # assign(paste0(’Simu’,i),data.frame(data))
+  fit1 <- vglm(cbind(y1,y2,y3) ~ x1+x2, family=sratio, data1)
+  theta1=coef(fit1)
+  fit2 <- vglm(cbind(y1,y2,y3) ~ x1, family=sratio, data1)
+  theta2=coef(fit2)
+ theta_matrix[i,1]=data.matrix(theta1)[1]
+ theta_matrix[i,2]=data.matrix(theta1)[3]
+ theta_matrix[i,3]=data.matrix(theta1)[5]
+ theta_matrix[i,4]=data.matrix(theta2)[2]
+ theta_matrix[i,5]=data.matrix(theta2)[4]
With the parameter matrix theta_matrix, the functions hfunc.temp and hprime.temp previously defined, we use EW_ForLion_MLM_Optimal() function to find a sample-based EW D-optimal approximate design as follows:
> set.seed(123)
> EW_forlion.design_MLM =EW_ForLion_MLM_Optimal(J=3 ,n.factor=c(0),
+ factor.level=list(c(0, 200)), hfunc=hfunc.temp, h.prime=hprime.temp,
+ bvec_matrix=theta_matrix, link="continuation", EW_Fi.func=EW_Fi_MLM_func,
+ delta0=1e-6, epsilon=1e-12, reltol=1e-6, delta=0.8, maxit=1000, random=TRUE,
+ nram=1, rowmax=NULL, Xini=NULL, random.initial=TRUE, nram.initial=1)
The sample-based EW D-optimal approximate design, denoted by $\boldsymbol{\xi}_{\rm SEW}$, is output below.
> print(EW_forlion.design_MLM)
Design Output
Count  X1        Allocation
1      148.7592  0.2598
2      149.5954  0.1394
3      103.4258  0.3979
4        0.0000  0.2030
m:
[1] 4
det:
[1] 58703238
convergence:
[1] TRUE
min.diff:
[1] 0.8363
x.close:
[1] 148.7592 149.5954
itmax:
[1] 13
For illustration purpose, we adopt $\delta=0.8$ as the merging threshold. The above Design Output shows that the reported EW D-optimal approximate design contains four design points, namely $0.0000$, $103.4258$, $148.7592$, $149.5954$, and their corresponding approximate allocations are $0.2030$, $0.3979$, $0.2598$, $0.1394$, respectively. The determinant of the expected Fisher information matrix is $|\hat{E}\{{\mathbf{F}}({\boldsymbol{\xi}}_{\rm SEW},\boldsymbol{\Theta})\}|=58,703,238$.
Obtaining an exact design from the sample-based EW D-optimal design ${\boldsymbol{\xi}}_{\rm SEW}$
Similarly to the locally D-optimal approximate design $\boldsymbol{\xi}$, in practice we need to convert the sample-based EW D-optimal approximate design ${\boldsymbol{\xi}}_{\rm SEW}$ into an exact design with prespecified grid level $L$ and the total number $N$ of experimental units. We may use the same function MLM_Exact_Design(). Instead of entering ForLion=TRUE and the parameter vector bvec, we need to set ForLion=FALSE indicating EW ForLion algorithm, and input the bootstrapped parameter matrix theta_matrix for argument bvec_matrix. The corresponding exact design with $L=0.1$ and $N=3,500$ is obtained by the following R code:
>  EW_ForLion_exact_L0.1<-MLM_Exact_Design(J=3, k.continuous=1,design_x=
+ EW_forlion.design_MLM\$x.factor,design_p=EW_forlion.design_MLM\$p,det.design=
+ EW_forlion.design_MLM\$det,p=5,ForLion=FALSE,bvec_matrix=theta_matrix,
+ delta2=1,L=0.1,N=3500,hfunc=hfunc.temp,link="continuation")
The obtained exact design, denoted by ${\boldsymbol{\xi}}^{\prime}_{\rm exact}$, is output as below:
> print(EW_ForLion_exact_L0.1)
Design Output
Count  X1        Allocation
1      103.4000  0.3979
2        0.0000  0.2030
3      149.1000  0.3991
ni.design:
[1] 1393  710 1397
rel.efficiency:
[1] 1.000056
Instead of four design points in ${\boldsymbol{\xi}}_{\rm SEW}$, the exact design ${\boldsymbol{\xi}}^{\prime}_{\rm exact}$ contains only three design points, namely $0,103.4$, and $149.1$, along with the integer-valued allocations $710,1393$, and $1397$. The relative efficiency of ${\boldsymbol{\xi}}^{\prime}_{\rm exact}$ with respect to ${\boldsymbol{\xi}}_{\rm SEW}$ is $(|\hat{E}\{{\mathbf{F}}({\boldsymbol{\xi}}^{\prime}_{\rm exact},\boldsymbol{\Theta})\}|/|\hat{E}\{{\mathbf{F}}({\boldsymbol{\xi}}_{\rm SEW},\boldsymbol{\Theta})\}|)^{1/p}=1.000056$ or $100.0056\%$.
4.2 A GLM example: Electrostatic discharge (ESD) experiment
Lukemire et al. (2019) revisited the electrostatic discharge (ESD) experiment, initially described by Whitman et al. (2006), as a GLM example (see also Huang et al. (2024) and Lin et al. (2025)). It involves a binary response, whether a certain part of the semiconductor fails, and five mixed experimental factors. The first four factors, namely LotA, LotB, ESD, and Pulse, take values in $\{-1,1\}$, while the fifth factor Voltage is continuous within the range $[25,45]$.
A logistic regression model is considered for this experiment:
$\text{logit}(\mu)=\beta_{0}+\beta_{1}\texttt{LotA}+\beta_{2}\texttt{LotB}+\beta_{3}\texttt{ESD}+\beta_{4}\texttt{Pulse}+\beta_{5}\texttt{Voltage}+\beta_{34}(\texttt{ESD}\times\texttt{Pulse})$
with parameters $\boldsymbol{\theta}=(\beta_{5},\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{34},\beta_{0})^{\top}$. Note that we let the coefficient $\beta_{5}$ for the continuous factor be the first, and the intercept be the last.
Finding a locally D-optimal approximate design
To use ForLion_GLM_Optimal() function, we first need to specify the function for generating the design matrix, and the assumed model parameters $\boldsymbol{\theta}=(0.35,1.50,-0.2,-0.15,0.25,0.4,-7.5)^{\top}$ adopted by Lukemire et al. (2019) and Huang et al. (2024):
> hfunc.temp = function(y) {c(y,y[4]*y[5],1);}; #design matrix
> beta.value = c(0.35,1.50,-0.2,-0.15,0.25,0.4,-7.5) #coefficient
> variable_names=c("Vol.","LotA","LotB","ESD","Pul.")
Next, we use the argument n.factor = c(0, 2, 2, 2, 2) to specify the experimental factor structure, where 0 represents the continuous factor Voltage (always come first) and the subsequent $2$’s represent binary discrete factors (four in total). The corresponding list of factor levels are provided in factor.level = list(c(25,45),c(-1,1),c(-1,1),c(-1,1),c(-1,1)), where c(25,45) stands for an interval, $[25,45]$, for the continuous factor, and c(-1,1) represents a set, $\{-1,1\}$, for a discrete factor. The corresponding R commands are listed as below:
> set.seed(482)
> forlion.glm=ForLion_GLM_Optimal(n.factor=c(0, 2, 2, 2, 2), factor.level=
+ list(c(25,45),c(-1,1),c(-1,1),c(-1,1),c(-1,1)), var_names=variable_names,
+ hfunc=hfunc.temp, bvec=beta.value, link="logit", reltol=1e-5, delta=0.1,
+ maxit=1000, random=FALSE, logscale=TRUE)
The results obtained from the GLM-adapted ForLion algorithm can be summarized and displayed using print() function. It provides a concise overview of the characteristics of the obtained optimal design, including the selected design points, their corresponding allocations, the determinant of the Fisher information matrix, etc.
> print(forlion.glm)
Design Output
Count  Vol.     LotA     LotB     ESD      Pul.     Allocation
1      25.0275  -1.0000   1.0000   1.0000  -1.0000  0.0432
2      25.1062  -1.0000   1.0000  -1.0000  -1.0000  0.0828
3      25.1957  -1.0000  -1.0000   1.0000  -1.0000  0.1100
4      28.5555  -1.0000  -1.0000  -1.0000   1.0000  0.0742
5      33.0930  -1.0000   1.0000   1.0000  -1.0000  0.0462
6      25.0000  -1.0000  -1.0000   1.0000   1.0000  0.0855
7      25.0000  -1.0000  -1.0000  -1.0000   1.0000  0.0339
8      29.1384  -1.0000   1.0000  -1.0000  -1.0000  0.0135
9      25.0000  -1.0000   1.0000   1.0000   1.0000  0.0923
10     25.0000   1.0000   1.0000   1.0000  -1.0000  0.1331
11     31.5543  -1.0000  -1.0000   1.0000  -1.0000  0.0018
12     25.0000   1.0000  -1.0000   1.0000  -1.0000  0.0136
13     25.0000  -1.0000   1.0000  -1.0000   1.0000  0.1013
14     25.0000  -1.0000  -1.0000  -1.0000  -1.0000  0.0865
15     32.8079  -1.0000   1.0000   1.0000  -1.0000  0.0822
m:
[1] 15
det:
[1] 1.256089e-05
convergence:
[1] TRUE
min.diff:
[1] 0.2851
x.close:
        [,1] [,2] [,3] [,4] [,5]
[1,] 33.0930   -1    1    1   -1
[2,] 32.8079   -1    1    1   -1
itmax:
[1] 49
The above Design Output shows that the obtained design, denoted by $\boldsymbol{\xi}$, contains $15$ design points in the table, whose levels are listed in the same order as the ones specified by factor.level. The last column of the table is the corresponding approximate allocations for the design points. The determinant of the Fisher information matrix, namely det, is $\left|\mathbf{F}(\boldsymbol{\xi},\boldsymbol{\theta})\right|=1.256089\times 10^{-5}$. In this locally D-optimal approximate design, the minimum Euclidean distance between the design points is equal to $0.2851$, and the closest pair of design points is reported by x.close. The algorithm converges (convergence:TRUE) with the number of iterations itmax=49.
Obtaining exact designs based on the locally D-optimal approximate design $\boldsymbol{\xi}$
The approximate design $\boldsymbol{\xi}$ specifies accurate levels of the continuous factor Voltage like $25.0275$, $25.1957$, $29.1384$, and $31.5543$. Those voltage levels may not be able to be maintained precisely in the experiment. If, for example, the voltage in this experiment can only be controlled to be a multiple of $L=0.1$, we may employ GLM_Exact_Design() function to convert $\boldsymbol{\xi}$ into a feasible exact design with modified voltage levels and integer-valued allocations. For illustration purpose, we set the total number of observations to $N=500$, the merging threshold delta2 to $0.5$, and the rounding level to $L=0.1$ for the only continuous factor Voltage. The corresponding exact design can be obtained by the following R code:
> ForLion_exactL0.1_ESD<-GLM_Exact_Design(k.continuous=1, design_x=
+ forlion.glm\$x.factor, design_p=forlion.glm\$p, var_names=variable_names,
+ det.design=forlion.glm\$det, p=7, ForLion=TRUE, bvec=beta.value, delta2=0.5,
+ L=0.1,N=500, hfunc=hfunc.temp, link="logit")
> print(ForLion_exactL0.1_ESD)
Design Output
Count  Vol.     LotA     LotB     ESD      Pul.     Allocation
1      25.0000  -1.0000   1.0000   1.0000  -1.0000  0.0432
2      25.1000  -1.0000   1.0000  -1.0000  -1.0000  0.0828
3      25.2000  -1.0000  -1.0000   1.0000  -1.0000  0.1100
4      28.6000  -1.0000  -1.0000  -1.0000   1.0000  0.0742
5      25.0000  -1.0000  -1.0000   1.0000   1.0000  0.0855
6      25.0000  -1.0000  -1.0000  -1.0000   1.0000  0.0339
7      29.1000  -1.0000   1.0000  -1.0000  -1.0000  0.0135
8      25.0000  -1.0000   1.0000   1.0000   1.0000  0.0923
9      25.0000   1.0000   1.0000   1.0000  -1.0000  0.1331
10     31.6000  -1.0000  -1.0000   1.0000  -1.0000  0.0018
11     25.0000   1.0000  -1.0000   1.0000  -1.0000  0.0136
12     25.0000  -1.0000   1.0000  -1.0000   1.0000  0.1013
13     25.0000  -1.0000  -1.0000  -1.0000  -1.0000  0.0865
14     32.9000  -1.0000   1.0000   1.0000  -1.0000  0.1284
ni.design:
 [1] 22 41 55 37 43 17  7 46 66  1  7 51 43 64
rel.efficiency:
[1] 1.000069
The Design Output shows that the obtained exact design, denoted by $\boldsymbol{\xi}_{\rm exact}$, only contains $14$ design points as listed in the table of design output. The levels of Voltage (listed as the first factor in the table) have been rounded to multiples of $L=0.1$. With $N=500$, the integer-values allocations are listed in ni.design. The relative efficiency of the exact design $\boldsymbol{\xi}_{\rm exact}$ with respect to the approximate design $\boldsymbol{\xi}$ is provided as rel.efficiency, that is, $(|{\mathbf{F}}(\boldsymbol{\xi}_{\rm exact},\boldsymbol{\theta})|/|{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})|)^{1/p}=1.000069$ or $100.0069\%$.
To illustrate how the exact design changes along with $N$, we generate two more exact designs with N=100, N=500, respectively, both with L=0.5. Both designs are listed in Table 5. Their relative efficiencies are $100.0529\%$ for N=100 and $100.1184\%$ for N=500.
Table 5: Exact designs with $L=0.5$ and different $N$’s for the ESD experiment
Support $N=100$ Support $N=500$
point Vol. LotA LotB ESD Pul. $n_{i}$ point Vol. LotA LotB ESD Pul. $n_{i}$
1 25 -1 1 1 -1 4 1 25 -1 1 1 -1 22
2 25 -1 1 -1 -1 8 2 25 -1 1 -1 -1 41
3 25 -1 -1 1 -1 11 3 25 -1 -1 1 -1 55
4 28.5 -1 -1 -1 1 8 4 28.5 -1 -1 -1 1 37
5 25 -1 -1 1 1 9 5 25 -1 -1 1 1 43
6 25 -1 -1 -1 1 4 6 25 -1 -1 -1 1 17
7 29 -1 1 -1 -1 1 7 29 -1 1 -1 -1 7
8 25 -1 1 1 1 9 8 25 -1 1 1 1 46
9 25 1 1 1 -1 13 9 25 1 1 1 -1 66
10 - - - - - - 10 31.5 -1 -1 1 -1 1
11 25 1 -1 1 -1 1 11 25 1 -1 1 -1 7
12 25 -1 1 -1 1 10 12 25 -1 1 -1 1 51
13 25 -1 -1 -1 -1 9 13 25 -1 -1 -1 -1 43
14 33 -1 1 1 -1 13 14 33 -1 1 1 -1 64
Finding integral-based EW D-optimal approximate design
To find a robust D-optimal approximate design against possibly misspecified $\boldsymbol{\theta}$, we adopt the prior distribution suggested by Huang et al. (2024). That is, we assume a prior distribution for $\boldsymbol{\Theta}$ (i.e., a randomized version of $\boldsymbol{\theta}$), such that, (i) all components of $\boldsymbol{\Theta}$ are independent of each other; and (ii) each component of $\boldsymbol{\Theta}$ follows a uniform distribution listed below:
$\beta_{0}\sim U(-8,-7),\quad\beta_{1}\sim U(1,2),\quad\beta_{2}\sim U(-0.3,-0.1),\quad\beta_{3}\sim U(-0.3,0),$
$\beta_{4}\sim U(0.1,0.4),\quad\beta_{5}\sim U(0.25,0.45),\quad\beta_{34}\sim U(0.35,0.45).$
To find an integral-based EW D-optimal approximate design given the prior distribution above, we first use two vectors, paras_lowerbound and paras_upperbound, to denote the lower bounds and upper bounds of the uniform distributions, respectively. Note that the orders of vector coordinates must match the same order as in factor.level. Then we can define the probability density function (pdf) of the prior distribution by gjoint function below:
> paras_lowerbound<-c(0.25,1,-0.3,-0.3,0.1,0.35,-8.0)
> paras_upperbound<-c(0.45,2,-0.1,0.0,0.4,0.45,-7.0)
> gjoint_b<- function(x) {
  Func_b<-1/(prod(paras_upperbound-paras_lowerbound))
  ##the prior distributions are follow uniform distribution
  return(Func_b)
By utilizing the EW_ForLion_GLM_Optimal() function with specified arguments, we obtain an integral-based EW D-optimal design as below.
> set.seed(482)
> EW_ForLion_ESD=EW_ForLion_GLM_Optimal(n.factor=c(0, 2, 2, 2, 2),  factor.level=
+ list(c(25,45),c(-1,1),c(-1,1),c(-1,1),c(-1,1)), var_names=variable_names,
+ hfunc=hfunc.temp, Integral_based=TRUE, joint_Func_b=gjoint_b, Lowerbounds=
+ paras_lowerbound, Upperbounds=paras_upperbound, link="logit", reltol=1e-6,
+ delta=0.01, maxit=1000, random=FALSE,  logscale=TRUE)
> print(EW_ForLion_ESD)
Design Output
Count  Vol.     LotA     LotB     ESD      Pul.     Allocation
1      25.0275  -1.0000   1.0000   1.0000  -1.0000  0.0756
2      25.1062  -1.0000   1.0000  -1.0000  -1.0000  0.0688
3      25.1210  -1.0000  -1.0000  -1.0000  -1.0000  0.0848
4      25.1297  -1.0000   1.0000  -1.0000   1.0000  0.0718
5      25.1957  -1.0000  -1.0000   1.0000  -1.0000  0.0585
6      25.2129  -1.0000   1.0000   1.0000   1.0000  0.0864
7      25.4221  -1.0000  -1.0000   1.0000   1.0000  0.0847
8      25.4990  -1.0000  -1.0000  -1.0000   1.0000  0.0645
9      25.0000   1.0000   1.0000  -1.0000   1.0000  0.0619
10     34.0891  -1.0000   1.0000  -1.0000  -1.0000  0.0158
11     39.0310  -1.0000   1.0000   1.0000  -1.0000  0.0932
12     25.0000   1.0000   1.0000  -1.0000  -1.0000  0.0362
13     25.0000   1.0000  -1.0000   1.0000  -1.0000  0.0745
14     37.2799  -1.0000  -1.0000   1.0000  -1.0000  0.0073
15     25.0000   1.0000   1.0000   1.0000   1.0000  0.0091
16     39.0638  -1.0000   1.0000   1.0000  -1.0000  0.0240
17     25.0000   1.0000   1.0000   1.0000  -1.0000  0.0439
18     33.5554  -1.0000  -1.0000  -1.0000   1.0000  0.0389
m:
[1] 18
det:
[1] 4.372488e-06
convergence:
[1] TRUE
min.diff:
[1] 0.0328
x.close:
        [,1] [,2] [,3] [,4] [,5]
[1,] 39.0310   -1    1    1   -1
[2,] 39.0638   -1    1    1   -1
itmax:
[1] 36
The reported integral-based EW D-optimal design, denoted by $\boldsymbol{\xi}_{\rm EW}$, consists of $18$ design points. The corresponding determinant det of the expected Fisher information matrix is $|E\{{\mathbf{F}}({\boldsymbol{\xi}}_{\rm EW},\boldsymbol{\Theta})\}|=4.372488\times 10^{-6}$. Having performed on a Windows 11 laptop with 32GB of RAM and a 13th Gen Intel Core i7-13700HX processor, with R version 4.4.2, the above procedure costs 2331 seconds.
Obtaining an exact design based on the integral-based EW D-optimal design $\boldsymbol{\xi}_{\rm EW}$
Similarly to the locally D-optimal design $\boldsymbol{\xi}$, we also need to convert the EW D-optimal approximate design $\boldsymbol{\xi}_{\rm EW}$ into an exact design for practical uses. The same function GLM_Exact_Design() can be applied, but with ForLion = FALSE indicating that the original design was obtained by an EW ForLion algorithm. In this case, we also need to input the pdf joint_Func_b for the prior distribution, along with the lower and upper bounds of ranges, namely Lowerbounds and Upperbounds. For illustration purpose, we still use the rounding level $L=0.1$ for the continuous factor and the total number of observations $N=500$.
> EW_ForLion_exactL0.1_ESD<-GLM_Exact_Design(k.continuous=1,design_x=
  + EW_ForLion_ESD\$x.factor,design_p=EW_ForLion_ESD\$p, var_names=variable_names,
  + det.design=EW_ForLion_ESD\$det,p=7,ForLion=FALSE,Integral_based=TRUE,
  + joint_Func_b=gjoint_b, Lowerbounds=paras_lowerbound, Upperbounds=
  + paras_upperbound,delta2 =0.5, L=0.1, N=500,hfunc=hfunc.temp,link="logit")
> print(EW_ForLion_exactL0.1_ESD)
Design Output
Count  Vol.     LotA     LotB     ESD      Pul.     Allocation
1      25.0000  -1.0000   1.0000   1.0000  -1.0000  0.0756
2      25.1000  -1.0000   1.0000  -1.0000  -1.0000  0.0688
3      25.1000  -1.0000  -1.0000  -1.0000  -1.0000  0.0848
4      25.1000  -1.0000   1.0000  -1.0000   1.0000  0.0718
5      25.2000  -1.0000  -1.0000   1.0000  -1.0000  0.0585
6      25.2000  -1.0000   1.0000   1.0000   1.0000  0.0864
7      25.4000  -1.0000  -1.0000   1.0000   1.0000  0.0847
8      25.5000  -1.0000  -1.0000  -1.0000   1.0000  0.0645
9      25.0000   1.0000   1.0000  -1.0000   1.0000  0.0619
10     34.1000  -1.0000   1.0000  -1.0000  -1.0000  0.0158
11     25.0000   1.0000   1.0000  -1.0000  -1.0000  0.0362
12     25.0000   1.0000  -1.0000   1.0000  -1.0000  0.0745
13     37.3000  -1.0000  -1.0000   1.0000  -1.0000  0.0073
14     25.0000   1.0000   1.0000   1.0000   1.0000  0.0091
15     25.0000   1.0000   1.0000   1.0000  -1.0000  0.0439
16     33.6000  -1.0000  -1.0000  -1.0000   1.0000  0.0389
17     39.0000  -1.0000   1.0000   1.0000  -1.0000  0.1173
ni.design:
 [1] 38 34 42 36 29 43 42 32 31  8 18 37  4  5 22 20 59
rel.efficiency:
[1] 1.000431
The reported exact design consists of $17$ design points, along with their corresponding allocations provided by ni.design. Its relative efficiency with respect to $\boldsymbol{\xi}_{\rm EW}$ is $1.000431$ or $100.0431\%$.
Finding a sample-based EW D-optimal approximate design
Given the same prior distribution for obtaining $\boldsymbol{\xi}_{\rm EW}$, we can also simulate $B=1,000$ random parameter vectors $\{\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{B}\}$ from the prior distribution, and then find a sample-based EW D-optimal approximate design based on the simulated parameter vectors.
> nrun = 1000
> set.seed(0713)
> b_0 = runif(nrun, -8, -7)
> b_1 = runif(nrun, 1, 2)
> b_2 = runif(nrun, -0.3, -0.1)
> b_3 = runif(nrun, -0.3, 0)
> b_4 = runif(nrun, 0.1, 0.4)
> b_5 = runif(nrun, 0.25, 0.45)
> b_34= runif(nrun, 0.35, 0.45)
> beta.matrix = cbind(b_5,b_1,b_2,b_3,b_4,b_34,b_0)
Similarly to obtaining $\boldsymbol{\xi}_{\rm EW}$, we can also apply the EW_ForLion_GLM_Optimal() function, but with Integral_based = FALSE indicating sample-based EW D-optimality, and input the sampled parameter matrix beta.matrix obtained above for argument b_matrix.
> set.seed(482)
> sample_EW_ForLion_ESD=EW_ForLion_GLM_Optimal(n.factor=c(0, 2, 2, 2, 2),
+ factor.level=list(c(25,45),c(-1,1),c(-1,1),c(-1,1),c(-1,1)), var_names=
+ variable_names, hfunc=hfunc.temp, Integral_based=FALSE,b_matrix=beta.matrix,
+ link="logit", reltol=1e-6, delta=0.01, maxit=1000, random=FALSE,logscale=TRUE)
> print(sample_EW_ForLion_ESD)
Design Output
Count  Vol.     LotA     LotB     ESD      Pul.     Allocation
1      25.0275  -1.0000   1.0000   1.0000  -1.0000  0.0753
2      25.1062  -1.0000   1.0000  -1.0000  -1.0000  0.0743
3      25.1210  -1.0000  -1.0000  -1.0000  -1.0000  0.0844
4      25.1297  -1.0000   1.0000  -1.0000   1.0000  0.0736
5      25.1957  -1.0000  -1.0000   1.0000  -1.0000  0.0642
6      25.2129  -1.0000   1.0000   1.0000   1.0000  0.0895
7      25.4221  -1.0000  -1.0000   1.0000   1.0000  0.0838
8      25.4990  -1.0000  -1.0000  -1.0000   1.0000  0.0659
9      38.6154  -1.0000   1.0000   1.0000  -1.0000  0.0953
10     33.6469  -1.0000   1.0000  -1.0000  -1.0000  0.0097
11     25.0000   1.0000   1.0000   1.0000   1.0000  0.0033
12     25.0000   1.0000   1.0000   1.0000  -1.0000  0.0495
13     25.0000   1.0000  -1.0000   1.0000  -1.0000  0.0705
14     36.8736  -1.0000  -1.0000   1.0000  -1.0000  0.0051
15     25.0000   1.0000   1.0000  -1.0000   1.0000  0.0612
16     33.2671  -1.0000  -1.0000  -1.0000   1.0000  0.0356
17     38.5979  -1.0000   1.0000   1.0000  -1.0000  0.0269
18     25.0000   1.0000   1.0000  -1.0000  -1.0000  0.0319
m:
[1] 18
det:
[1] 4.038136e-06
convergence:
[1] TRUE
min.diff:
[1] 0.0175
x.close:
        [,1] [,2] [,3] [,4] [,5]
[1,] 38.6154   -1    1    1   -1
[2,] 38.5979   -1    1    1   -1
itmax:
[1] 769
The obtained sample-based EW D-optimal approximate design, denoted by $\boldsymbol{\xi}_{\rm SEW}$, contains m $=18$ design points, with det $=$ $|\hat{E}\{{\mathbf{F}}({\boldsymbol{\xi}}_{\rm SEW},\boldsymbol{\Theta})\}|=4.038136\times 10^{-6}$.
Comparing sample-based and integral-based EW D-optimal designs
According to a simulation study done by Lin et al. (2025) on a minimizing surface defects experiment, sample-based EW D-optimal designs are fairly robust in terms of relative efficiencies against different set of simulated parameter vectors.
In this paper, we use this ESD experiment to compare the integral-based EW D-optimal design $\boldsymbol{\xi}_{\rm EW}$ with sample-based EW D-optimal designs based on six different sets of simulated parameter vectors. More specifically, for $b=1$ (representing $B=100$) and $b=2$ (representing $B=1,000$), we simulate three random sets of parameter vectors of size $B$, labeled by $j=1,2,3$, and find the corresponding sample-based EW D-optimal designs, denoted by $\boldsymbol{\xi}_{bj}$ Then we calculate the relative efficiencies of $\boldsymbol{\xi}_{bj}$ with respect to $\boldsymbol{\xi}_{\rm EW}$, in terms of the integral-based EW D-optimality. That is,
$\left(\frac{|E\{{\mathbf{F}}(\boldsymbol{\xi}_{bj},\boldsymbol{\Theta})\}|}{|E\{{\mathbf{F}}(\boldsymbol{\xi}_{\rm EW},\boldsymbol{\Theta})\}|}\right)^{1/p}\ ,$
for $b=1,2$, $j=1,2,3$, and $p=7$. The relative efficiencies are shown in the following matrix,
[…]
with the minimum relative efficiency $0.9923964$ or $99.23964\%$, which is fairly high.
For readers’ reference, the numbers of distinct design points contained in $\boldsymbol{\xi}_{bj}$’s are listed below,
[…]
which vary from design to design, but not so different from each other.
5 Summary and discussion
In this paper, we introduce the ForLion package, which facilitates the users to find D-optimal and EW D-optimal design of experiments involving discrete factors only, continuous factors only, or mixed factors. It utilizes the ForLion and EW ForLion algorithms, and conducts an exhaustive search within the design space, ensuring the identification of the D-optimal designs while reducing the numbers of distinct experimental settings significantly.
The current version 0.3.0 of ForLion package supports both GLM and MLM for various experimental scenarios. It is worth noting that a regular linear regression model is a special case of GLM with an identity link function, and covered by ForLion package as well. In this package, the functions ForLion_MLM_Optimal() and ForLion_GLM_Optimal() can be used for determining locally D-optimal approximate designs, if the experimenter is certain about the values of the parameters. When the true parameter values are unknown, while either a prior distribution $Q(\cdot)$ on $\boldsymbol{\Theta}$ or a dataset from a pilot study is available, we provide functions EW_ForLion_MLM_Optimal() and EW_ForLion_GLM_Optimal() to find EW D-optimal approximate designs. Having obtained D-optimal approximate designs, we also provide functions GLM_Exact_Design() for GLM and MLM_Exact_Design() for MLM to convert the approximate designs with values of possible continuous factors into exact designs with user-specified grid levels and the total number of experimental units. By using this rounding algorithm, the yielded exact designs can maintain high relative efficiency with respect to the D-optimal approximate designs, and may further reduce the number of distinct experimental settings.
Following Huang et al. (2024) and Lin et al. (2025), the current ForLion package concentrates on D-optimality, which is not only the most commonly used criterion in optimal design theory, but often leads to a design that performs well with respect to other criteria. Nevertheless, it can be extended to other criteria, given a corresponding lift-one algorithm being developed.